The graph to the right shows a region of feasible solutions. Use this region to find maximum and minimum values of the given objective functions, and the locations of these values on the graph. 10 (2, 8) (a) z 6x+9y (b)2-x + 3y (7, 5) (0 어 (a) What is the maximum of z 6x + 9y? Select the correct answer below and, if necessary, fill in the answer boxes to complete your choice. A. The maximum value of the objective function Z 6x+9y is 87 and is located at the point 0- 10 (7,5) Tyne an evact answerWhat is the minimum of zx+3y? Select the correct answer below and, if necessary, fill in the answer boxes to complete your choice. O A. The minimum value of the objective function zx+ 3y is and is located at (Type an exact answer.) O B. The minimum does not exist.
Answer #1

Answer b.

Z = x + 3y

At A (0, 6):

Z = 0 + 3 * 6
Z = 18

At B (3, 1):

Z = 3 + 3 * 1
Z = 6

At C (7, 5):

Z = 7 + 3 * 5
Z = 22

At D (2, 8):

Z = 2 + 3 * 8
Z = 26

The minimum value of the objective function z = x + 3y is 6 and is located at (3, 1)